# Granary: Speech Recognition and Translation Dataset in 25 European Languages

### v1

@@ 1 Introduction
Efforts to open-source speech data remain limited. Notable examples include YODAS [4] and YouTube-Commons (YTC) [5], which provide large-scale datasets with labels derived from YouTube captions, albeit without guarantees regarding quality or source reliability. More recently, MOSEL [6] has released pseudo-generated labels for European languages, covering datasets such as VoxPopuli [7] and LibriLight [8]. However, MOSEL’s pseudo-labeling pipeline has not been publicly released, limiting its applicability to other datasets. Other community efforts have highlighted corpus creation pipelines, but these remain restricted to human-generated data and cover only a limited number of languages [9].
@@
  • Open-source large-scale speech processing pipeline
  • Efficient method for generating translation pairs from ASR transcripts across 25 languages
  • 643k hours of high-quality pseudo-labeled data for 25 languages
  • Evaluation of the quality of pseudo-labeled data against the MOSEL pipeline for both high- and low-resource languages
@@ 3.2.2 LLM Inference
We perform LLM inference on processed ASR data using the translation prompt from EuroLLM’s model card. For optimal speed, we use greedy inference with vLLM. We also experimented with beam search, which provided a slight improvement for EuroLLM 1.7B but had diminishing returns for the 9B model, making the added computational cost unjustifiable.
@@ 3.2.3 Filtration
Our data filtration pipeline is implemented in NeMo-Curator [^3] https://github.com/NVIDIA/NeMo-Curator/, a GPU-accelerated data curation toolkit. Our filtration steps include a re-implementation of the length ratio filtering step from Moses [^4] https://github.com/moses-smt/mosesdecoder/blob/master/scripts/training/clean-corpus-n.perl, character histogram filtering [23], FastText language ID [24], as well as Quality Estimation filtration [22]. For the last step, we use cometoid-wmt23 model [25] through PyMarian interface [26] as our quality estimation setup. The resulting pipeline is very efficient and scalable. For instance, when scaling over 8 nodes with 8 NVIDIA A100 GPUs each, filtering MOSEL dataset took only 47 minutes.

### v2

@@ 1 Introduction
Efforts to open-source speech data remain limited. Notable examples include YODAS [4] and YouTube-Commons (YTC) [5], which provide large-scale datasets with labels derived from YouTube captions, albeit without guarantees regarding quality or source reliability. More recently, MOSEL [6] has released pseudo-generated labels for European languages, covering datasets such as VoxPopuli [7] and LibriLight [8]. Other community efforts have highlighted corpus creation pipelines, but these remain restricted to human-generated data and cover only a limited number of languages [9].
@@ 3.2.2 LLM Inference
We perform LLM inference on processed ASR data using the translation prompt from EuroLLM’s model card. For optimal speed, we use greedy inference with vLLM. We also experimented with beam search, which provided a slight improvement for EuroLLM-1.7B but had diminishing returns for the 9B model, making the added computational cost unjustifiable.
@@ 3.2.3 Filtration
Our data filtration pipeline is implemented in NeMo-Curator [^4] https://github.com/NVIDIA/NeMo-Curator/, a GPU-accelerated data curation toolkit. Our filtration steps include a re-implementation of the length ratio filtering step from Moses  [^5] https://github.com/moses-smt/.../clean-corpus-n.perl , character histogram filtering [23], FastText language ID [24], as well as Quality Estimation filtration [22]. For the last step, we use cometoid-wmt23 model [25] through PyMarian interface [26] as our quality estimation setup. The resulting pipeline is very efficient and scalable. For instance, when scaling over 8 nodes with 8 NVIDIA A100 GPUs each, filtering MOSEL dataset took only 47 minutes.